A sample of radioactive material contains 1.00 x 10^15 atoms and has an activity of 6.00 x 10^11 Bq. What is it's half-life?
1
Expert's answer
2021-02-17T14:43:11-0500

The relationship between the activity "R" and the number of nuclei in the sample, "N", can be written as follows:


"R=N\lambda,"

here, "\lambda" is decay constant.

Then, from this formula we can find the decay constant:


"\lambda=\dfrac{R}{N}."

The relationship between the half-life of a sample, "T_{1\/2}", and its decay constant "\lambda" can be written as follows:


"T_{1/2}=\dfrac{ln2}{\lambda}."

Substituting "\lambda" into the previous formula, we get:


"T_{1/2}=\dfrac{Rln2}{N}=\dfrac{1.0\cdot10^{15}\cdot ln2}{6.0\cdot10^{11}\ Bq}=1.16\cdot10^3\ s=19.3\ min."